# Stabilization of intersection Betti numbers for moduli spaces of one-dimensional sheaves on surfaces

### v1

@@ 2.1. Relative Hilbert schemes of points
The morphism $\pi^{[n]}:{\mathcal{C}}^{[n]}\to|\beta|$ has fibers of the same dimension $n$.
There is a natural morphism $\sigma_{n}:{\mathcal{C}}^{[n+1]}\to S^{[n+1]}$ defined by
(2.3) $Z\subset C]\mapsto[Z\subset S],$
where $C\in|\beta|$ and $Z\subset C$ is a 0-dimensional closed subscheme of length $n+1$. The fiber of $\sigma_{n}$ over $Z\subset S]\in S^{[n+1]}$ is the projective space $\mathbb{P}(H^{0}(S,{\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta))^{\vee})$, where ${\mathcal{I}}_{Z}$ is the ideal sheaf of $Z\subset S$.
2.2. Positivity conditions
@@
We recall the following notion of $k$-very ampleness, which generalizes base-point-freeness (0-very ampleness) and very ampleness ($1$-very ampleness).
@@
An effective divisor $\beta$ on $S$ is called $k$-very ample $(k\in\mathbb{Z}_{\geq 0})$ if for any 0-dimensional closed subscheme $Z\subset S$ of length $k+1$, the restriction map
@@
The definition immediately implies two properties, which will be used in the sequel:
@@
    If $\beta$ is ample, then for any $k\in\mathbb{Z}_{\geq 0}$, there exists $d_{k}\in\mathbb{Z}_{>0}$ such that the multiple $d\beta$ is $k$-very ample for every integer $d\geq d_{k}$.
A well-known connection between $k$-very ampleness and the structure of ${\mathcal{C}}^{[k+1]}$ as a scheme over $S^{[k+1]}$ via $\sigma_{k}$ is as follows (e.g., [34, Proposition 2.6]).
Proposition 2.3.
2.3. Perverse filtrations
Next we recall some generalities of perverse filtrations. Let $\phi:X\to Y$ be a proper morphism between smooth quasiprojective varieties. We assume that $\dim X=a$, $\dim Y=b$, and all fibers of $\phi$ have dimension aba-b. The perverse filtration
$P_{0}H^{m}(X)\subset P_{1}H^{m}(X)\subset\dots\subset H^{m}(X)$
induced by $\phi$ is an increasing filtration on the rational cohomology $H^{*}(X)$, governed by the topology of the morphism $\phi$. It is defined to be
$P_{k}H^{m}(X)\vcentcolon=\mathrm{Im}\left\{H^{m-b}(Y,{{}^{\mathbf{p}}\tau_{\leq k}}(R\phi_{*}{\mathbb{Q}}_{X}[b]))\to H^{m-b}(Y,R\phi_{*}{\mathbb{Q}}_{X}[b])\right\}\subset H^{m}(X),$
where $^{\mathbf{p}}\tau_{\leq\bullet}$ is the perverse truncation functor [2]. If we apply the decomposition theorem to $\phi:X\to Y$, we obtain
$R\phi_{*}{\mathbb{Q}}_{X}[b]\simeq\bigoplus_{i=0}^{2a-2b}\mathcal{P}_{i}[-i]\in D^{b}_{c}(Y)$
with $\mathcal{P}_{i}$ a semisimple perverse sheaf on $Y$. There is an identity
$P_{k}H^{m}(X)=\mathrm{Im}\Big\{H^{m-b}(Y,\bigoplus_{i=0}^{k}\mathcal{P}_{i}[-i])\to H^{m}(X)\Big\}.$
The dimension of the graded piece $\mathrm{Gr}_{i}^{P}H^{i+j}(X)\vcentcolon=P_{i}H^{i+j}(X)/P_{i-1}H^{i+j}(X)$ is denoted by
(2.4) $n_{\phi}^{i,j}\vcentcolon=\dim_{\mathbb{Q}}\mathrm{Gr}_{i}^{P}H^{i+j}(X)=\dim_{\mathbb{Q}}H^{j-b}(Y,{\mathcal{P}}_{i}).$
Note that for the Hilbert–Chow morphism $h:M_{\beta,\chi}\to|\beta|$, the fibers of $h$ over integral curves in $|\beta|$ are the compactified Jacobians of degree $p_{a}(\beta)-1+\chi$ and have the same dimension $p_{a}(\beta)$. Later, we will use the following result of Maulik–Yun [22, Theorem 2.13] and Migliorini–Shende [24, Corollary 2].
Theorem 2.4 ([22, 24] [^1] The result of [22, 24] can be applied here since the smoothness of $h^{-1}(U)$ implies that all relative compactified Jacobians $\overline{J}_{U}^{e}$ ($e\in\mathbb{Z}$) are smooth by the existence of a section of the universal curve (étale locally on $U$, avoiding the singularities of fibers) and so are all relative Hilbert schemes ${\mathcal{C}}^{[\ell]}_{U}$ ($\ell\in\mathbb{Z}_{\geq 0}$) by [32, Proposition 14].).
Suppose that $|\beta|$ contains a connected smooth curve. Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth, and let $P_{\bullet}H^{*}(h^{-1}(U))$ be the perverse filtration induced by the restriction $h_{U}:h^{-1}(U)\to U$ of $h$. For any $\ell,m\in\mathbb{Z}_{\geq 0}$, we have an isomorphism of vector spaces
(2.5) $H^{m}\left({\mathcal{C}}^{[\ell]}_{U}\right)\cong\sum_{i+j\leq\ell;\,i,j\geq 0}\mathrm{Gr}_{i}^{P}H^{m-2j}\left(h^{-1}(U)\right).$
Remark 2.5.
When $\ell=0$, the isomorphism (2.5) reads
$H^{m}(U)\cong\mathrm{Gr}_{0}^{P}H^{m}\left(h^{-1}(U)\right).$
This isomorphism can be obtained as a consequence of Ngô’s support theorem [26, Théorème 7.2.1]. This theorem shows that $h_{U}:h^{-1}(U)\to U$ has full support; see [22, Theorem 2.4, §2.5] or [24, Corollary 9]. More precisely, there is an isomorphism
$Rh_{U*}{\mathbb{Q}}_{h^{-1}(U)}[\dim U]\cong\bigoplus_{i=0}^{2p_{a}(\beta)}\operatorname{IC}(\wedge^{i}R^{1}\pi^{s}_{*}{\mathbb{Q}}_{{\mathcal{C}}^{\mathrm{sm}}})[-i],$
where $\pi^{s}:{\mathcal{C}}^{\mathrm{sm}}\to|\beta|^{\mathrm{sm}}$ is the universal family of smooth curves in $|\beta|$.
@@ 3. Singular locus of relative compactified Jacobians
The goal of this section is to prove Theorem 1.2 about the smoothness of a big open subscheme of $h^{-1}(|\beta|^{\mathrm{int}})$. We begin by reviewing some basic notions of plane curve singularities and deformation theory.
3.1. Singularities and deformations of planar curves
@@ 3.1.1. Singularity invariants
We first recall some basic invariants of isolated singularities in planar curves.
@@
$\hat{\mathcal{O}}_{C,p}\cong\mathbb{C}[[x,y]]/(f).$
Thus, the Milnor and Tjurina numbers, and the $\delta$-invariant of $C$ at $p$ can be defined as those of $f$ in Definition 3.1, denoted by $\mu_{p}(C)$, $\tau_{p}(C)$, and $\delta_{p}(C)$, respectively.
For completeness, we give a proof of the following standard interpretation of the total Tjurina number of $C$ in terms of the Ext sheaf ${\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})$, where $\Omega_{C}$ is the cotangent sheaf of $C$.
@@
The sheaf $\Omega_{C}$ is locally free over the smooth locus $C\setminus C_{\mathrm{sing}}$ of $C$, so ${\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})$ is a torsion sheaf supported on $C_{\mathrm{sing}}$. Let $p\in C_{\mathrm{sing}}$ and assume that the singularity of $C$ at $p$ is analytically isomorphic to that of $\{f=0\}$ at $(0,0)$ for $f\in(x,y)\subset\mathbb{C}[x,y]$ without multiple factors. The lemma can be proved locally around $p$. For $R\vcentcolon=\mathbb{C}[x,y]/(f)$, we have $\Omega_{R}=(Rdx\oplus Rdy)/(\partial_{x}fdx+\partial_{y}fdy)$. By the conormal exact sequence
@@
we conclude that $\operatorname{Ext}^{1}(\Omega_{R},R)=R/(\partial_{x}f,\partial_{y}f)$, from which the result follows. ∎
@@ 3.1.2. Smoothness from local versality
Let $\pi_{B}:{\mathcal{C}}_{B}\to B$ be a family of reduced planar curves, i.e., a proper and flat morphism whose fibers are reduced curves with at worst planar singularities. Assume that $B$ is a smooth variety. There is a natural map, called the local Kodaira–Spencer map:
(3.7) $k_{\pi_{B},b}^{\text{loc}}:T_{b}B\to\prod_{p\in({\mathcal{C}}_{b})_{\mathrm{sing}}}T\mathrm{Def}_{{\mathcal{C}}_{b},p}=H^{0}({\mathcal{C}}_{b},{\mathcal{E}xt}^{1}(\Omega_{{\mathcal{C}}_{b}},{\mathcal{O}}_{{\mathcal{C}}_{b}})),$
where $({\mathcal{C}}_{b})_{\mathrm{sing}}$ is the singular locus of ${\mathcal{C}}_{b}\vcentcolon=\pi_{B}^{-1}(b)$, and $T\mathrm{Def}_{{\mathcal{C}}_{b},p}$ is the tangent space of the deformation functor of the complete local $\mathbb{C}$-algebra $\hat{{\mathcal{O}}}_{{\mathcal{C}}_{b},p}$.
@@
The family $\pi_{B}:{\mathcal{C}}_{B}\to B$ is called locally versal if for any point $b\in B$, the local Kodaira–Spencer map $k_{\pi_{B},b}^{\text{loc}}$ in (3.7) is surjective.
If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a family of integral planar curves, then for any $e\in\mathbb{Z}$, denote by
$p_{B,e}:\overline{J}_{B}^{e}\to B$
the relative compactified Jacobian of degree $e$ associated with $\pi_{B}$, whose fiber over a point $b\in B$ is the moduli space of rank $1$, torsion-free sheaves of degree $e$ (or equivalently, Euler characteristic $e+1-p_{a}({\mathcal{C}}_{b})$) on ${\mathcal{C}}_{b}$. Since $\pi_{B}$ admits, étale locally on $B$, a section which avoids the singularities of fibers of $\pi_{B}$, the smoothness of $\overline{J}_{B}^{e}$ for any $e$ is equivalent to the smoothness of $\overline{J}_{B}^{0}$. Hence, the following result follows directly from [7, Corollary B.3].
Theorem 3.5 ([7]).
If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a locally versal family of integral planar curves, then $\overline{J}_{B}^{e}$ is smooth for any $e\in\mathbb{Z}$.

### v2

@@ 2.1. Moduli spaces of one-dimensional sheaves
Suppose that $K_{S}$ is nef and $|\beta|$ contains a smooth connected curve. If $\dim|\beta|=\chi({\mathcal{O}}_{S}(\beta))-1$, then
  1. (i)
    the moduli space $M_{\beta,\chi}$ is irreducible;
  2. (ii)
    all fibers of $h:M_{\beta,\chi}\to|\beta|$ have dimension $p_{a}(\beta)$.
Proof.
(i) It suffices to show that $Q_{\beta}$ is irreducible. We use the deformation theory of sheaves and quotients (see [9, Theorem 6.4.9, Proposition 6.5.1] and [16, 2.A.8, Theorem 4.5.3]). For any $\xi\in Q_{\beta}$ representing the short exact sequence
(2.4) $0\to{\mathcal{K}}\to{\mathcal{O}}_{S}(-nH)^{\oplus N}\to{\mathcal{F}}\to 0,$
the dimension of $Q_{\beta}$ at $\xi$, denoted by $\dim_{\xi}Q_{\beta}$, satisfies (cf. [16, Proposition 4.5.9])
(2.5) $\dim_{\xi}Q_{\beta}$ $\geq\dim_{\xi}Q-\dim\operatorname{Pic}(S)$
$\geq\operatorname{ext}^{0}({\mathcal{K}},{\mathcal{F}})-(\operatorname{ext}^{2}({\mathcal{F}},{\mathcal{F}})-h^{2}({\mathcal{O}}_{S}))-h^{1}({\mathcal{O}}_{S})$
$=N^{2}+\beta^{2}+\chi({\mathcal{O}}_{S})-1,$
where the last equality follows from the long exact sequence of Ext groups for (2.4) and the Riemann–Roch formula. By [42, Proposition 2.2], the open subscheme $h^{-1}(|\beta|^{\mathrm{int}})$ is irreducible and contains the smooth open subscheme $h^{-1}(|\beta|^{\mathrm{sm}})$, where $|\beta|^{\mathrm{sm}}\subset|\beta|$ is the open subscheme parametrizing smooth curves. Then by [16, Corollary 4.3.5], the open subscheme $Q_{\beta}^{\mathrm{int}}\subset Q_{\beta}$ parametrizing those quotient sheaves with integral support is irreducible and contains the smooth open subscheme $Q_{\beta}^{\mathrm{sm}}$ parametrizing those quotient sheaves with smooth support. There is a unique irreducible component $Y\subset Q_{\beta}$ containing $Q_{\beta}^{\mathrm{int}}$. If $Y\not=Q_{\beta}$, by taking an irreducible component $Y^{\prime}\neq Y$ and noting that $Y^{\prime}\cap Q_{\beta}^{\mathrm{sm}}=\varnothing$, we deduce from Lemma 2.1 and [43, Tag 0DS4] that
$\dim Y^{\prime}<\dim\operatorname{GL}(N)+\dim|\beta|+(p_{a}(\beta)-1)=N^{2}+\beta^{2}+\chi({\mathcal{O}}_{S})-1,$
which contradicts (2.5). Hence, $Q_{\beta}=Y$ is irreducible, and so is $M_{\beta,\chi}$.
(ii) By the upper semicontinuity of the fiber dimension of $h$, $\dim h^{-1}(C)\geq p_{a}(\beta)$ for any curve $C\in|\beta|$. Note that ${\mathcal{M}}_{\beta,\chi}$ is also irreducible since $Q_{\beta}$ is. Consider the morphism $\nu:{\mathcal{M}}_{\beta,\chi}\to M_{\beta,\chi}$ and its restriction $\nu|_{\tilde{h}^{-1}(C)}:\tilde{h}^{-1}(C)\to h^{-1}(C)$. By Lemma 2.1 and [43, Tag 0DS4], we have
$\dim\tilde{h}^{-1}(C)-\dim h^{-1}(C)=(p_{a}(\beta)-1)-\dim h^{-1}(C)\geq-1=\dim{\mathcal{M}}_{\beta,\chi}-\dim M_{\beta,\chi},$
which forces $\dim h^{-1}(C)=p_{a}(\beta)$. ∎
Remark 2.3.
If $\beta$ is an ample divisor on a K3 (resp. an Abelian) surface $S$ satisfying $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)\geq 2$, Proposition 2.2 (ii) can also be obtained from the symplectic geometry of $M_{\beta,\chi}$: [^1] The authors are grateful to Mirko Mauri for pointing out this alternative approach and the relevant references via private communication. Using [5, Propositions 3.5, 3.14], one can show that $M_{\beta,\chi}$ (resp. $K_{\beta,\chi}$) is a primitive symplectic variety (e.g., [5, Definition 3.3]) admitting a Lagrangian fibration over $|\beta|$ (with no restrictions on $H$ or $\chi$), and thus all fibers of $h:M_{\beta,\chi}\to|\beta|$ have dimension $p_{a}(\beta)$ by [38, Theorem 3]. Here, when $S$ is an Abelian surface with identity $0_{S}$, we denote by $K_{\beta,\chi}$ the fiber $\widehat{\varphi}^{\,-1}(0_{S})$ of the isotrivial fibration $\widehat{\varphi}:M_{\beta,\chi}\to S$ (cf. [45, Definition 4.1]).
2.2. Intersection cohomology and perverse filtrations
For a quasi-projective variety $X$, let $\operatorname{IC}_{X}$ denote the intersection cohomology complex of $X$ with $\mathbb{Q}$}-coefficients and middle perversity. The intersection cohomology groups of $X$ are defined as the hypercohomology groups of $\operatorname{IC}_{X}$:
$I\!H^{k}(X)\vcentcolon={\mathbb{H}}^{k-\dim X}(X,\operatorname{IC}_{X}).$
Note that $I\!H^{k}(X)$ can also be defined if $X$ is allowed to be non-reduced (by the topological invariance of intersection cohomology) or reducible (even not pure dimensional, as shown in [6, §4.6]). When $X$ is smooth, $I\!H^{k}(X)$ can be identified with the rational cohomology group $H^{k}(X)$.
Let $\varphi:X\to Y$ be a proper morphism between quasi-projective varieties. We assume that $\dim X=a$, $\dim Y=b$, and all fibers of $\varphi$ have dimension aba-b. The perverse filtration
$P_{0}I\!H^{k}(X)\subset P_{1}I\!H^{k}(X)\subset\dots\subset I\!H^{k}(X)$
induced by $\varphi$ is an increasing filtration on $I\!H^{k}(X)$, governed by the topology of the morphism $\varphi$. It is defined by
$P_{i}I\!H^{k}(X)\vcentcolon=\mathrm{Im}\left\{{\mathbb{H}}^{k-b}(Y,{{}^{\mathbf{p}}\tau_{\leq i}}(R\varphi_{*}\operatorname{IC}_{X}[b-a]))\to{\mathbb{H}}^{k-b}(Y,R\varphi_{*}\operatorname{IC}_{X}[b-a])\right\},$
where $^{\mathbf{p}}\tau_{\leq\bullet}$ is the perverse truncation functor. We denote the dimension of the graded piece $\mathrm{Gr}_{i}^{P}I\!H^{i+j}(X)\vcentcolon=P_{i}I\!H^{i+j}(X)/P_{i-1}I\!H^{i+j}(X)$ by
(2.6) $n_{\varphi}^{i,j}\vcentcolon=\dim_{\mathbb{Q}}\mathrm{Gr}_{i}^{P}I\!H^{i+j}(X).$
When $X$ is smooth, the decomposition theorem [2] applied to $\varphi:X\to Y$ yields
$R\varphi_{*}{\mathbb{Q}}_{X}[b]\simeq\bigoplus_{i=0}^{2a-2b}\mathcal{P}_{i}[-i],$
where $\mathcal{P}_{i}$ are semisimple perverse sheaves on $Y$; in this case, $n_{\varphi}^{i,j}=\dim_{\mathbb{Q}}{\mathbb{H}}^{j-b}(Y,{\mathcal{P}}_{i})$.
For our purposes, we mainly consider the Hilbert–Chow morphism $h:M_{\beta,\chi}\to|\beta|$ and its restrictions. The invariants $n_{\varphi}^{i,j}$ with $\varphi$ taken to be (the restrictions of) $h$ will play a central role in Section 4.
2.3. Positivity conditions for divisors
@@
To study the asymptotic behavior of $M_{\beta,\chi}$, we recall the definition of $k$-very ampleness for divisors on $S$, which generalizes base-point-freeness (0-very ampleness) and very ampleness ($1$-very ampleness).
@@
An effective divisor $\beta$ on $S$ is called $k$-very ample if for any 0-dimensional closed subscheme $Z\subset S$ of length $k+1$, the restriction map
@@
When $h^{1}({\mathcal{O}}_{S}(\beta))=0$, the restriction map $r_{Z}$ is surjective if and only if $h^{1}({\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta))=0$ by the long exact sequence of cohomology groups for the short exact sequence
$0\to{\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta)\to{\mathcal{O}}_{S}(\beta)\to{\mathcal{O}}_{S}(\beta)|_{Z}\to 0,$
where ${\mathcal{I}}_{Z}$ is the ideal sheaf of $Z\subset S$. Definition 2.4 immediately implies two properties, which will be used in the subsequent sections:
@@
    If $\beta$ is ample, then for any $k\in\mathbb{Z}_{\geq 0}$, there exists $d_{k}\in\mathbb{Z}_{>0}$ such that $d\beta$ is $k$-very ample for every integer $d\geq d_{k}$.
@@ 3. Singular locus of relative compactified Jacobians
The goal of this section is to prove Theorem 1.2. We begin by reviewing some basic notions of plane curve singularities and deformation theory.
3.1. Singularities and deformations of locally planar curves
@@ 3.1.1. Singularity invariants
We first recall some analytic invariants of isolated singularities in locally planar curves.
@@
$\widehat{\mathcal{O}}_{C,p}\cong\mathbb{C}[[x,y]]/(f).$
Thus, the Milnor number, the Tjurina number, and the $\delta$-invariant of $C$ at $p$ can be defined as those of $f$ in Definition 3.1, denoted by $\mu_{p}(C)$, $\tau_{p}(C)$, and $\delta_{p}(C)$, respectively.
For completeness, we give a proof of the following standard interpretation (cf. [13, Ch. II, Exercise 1.4.2]) of the total Tjurina number of $C$ in terms of the Ext sheaf ${\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})$, where $\Omega_{C}$ is the cotangent sheaf of $C$.
@@
The sheaf $\Omega_{C}$ is locally free over the smooth locus $C\setminus C_{\mathrm{sing}}$ of $C$, so ${\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})$ is a torsion sheaf supported on $C_{\mathrm{sing}}$. Let $p\in C_{\mathrm{sing}}$ and assume that the singularity of $C$ at $p$ is analytically isomorphic to that of $\{f=0\}$ at $(0,0)$ for $f\in(x,y)\subset\mathbb{C}[x,y]$ without multiple factors. The lemma can be proved locally around $p$. For $R\vcentcolon=\mathbb{C}[x,y]/(f)$, we have $\Omega_{R}\cong(Rdx\oplus Rdy)/(\partial_{x}fdx+\partial_{y}fdy)$. Using the conormal exact sequence
@@
we obtain $\operatorname{Ext}^{1}(\Omega_{R},R)\cong R/(\partial_{x}f,\partial_{y}f)$, from which the result follows. ∎
@@ 3.1.2. Smoothness from local versality
Let $\pi_{B}:{\mathcal{C}}_{B}\to B$ be a family of reduced, locally planar curves, i.e., a proper and flat morphism whose fibers are reduced curves with at worst planar singularities. Assume that $B$ is a smooth variety. There is a natural map (cf.  [13, Ch. II, Lemma 1.20, Exercise 1.4.4])
(3.8) $\kappa_{\pi_{B},b}^{\text{loc}}:T_{b}B\to\prod_{p\in({\mathcal{C}}_{b})_{\mathrm{sing}}}T\mathrm{Def}_{{\mathcal{C}}_{b},p}=H^{0}({\mathcal{C}}_{b},{\mathcal{E}xt}^{1}(\Omega_{{\mathcal{C}}_{b}},{\mathcal{O}}_{{\mathcal{C}}_{b}})),$
called the local Kodaira–Spencer map, where $({\mathcal{C}}_{b})_{\mathrm{sing}}$ is the singular locus of ${\mathcal{C}}_{b}\vcentcolon=\pi_{B}^{-1}(b)$, and $T\mathrm{Def}_{{\mathcal{C}}_{b},p}$ is the tangent space of the deformation functor of the complete local $\mathbb{C}$-algebra $\widehat{{\mathcal{O}}}_{{\mathcal{C}}_{b},p}$.
@@
The family $\pi_{B}:{\mathcal{C}}_{B}\to B$ is called locally versal if for any point $b\in B$, the local Kodaira–Spencer map $\kappa_{\pi_{B},b}^{\text{loc}}$ in (3.8) is surjective.
If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a family of integral, locally planar curves, then for any $e\in\mathbb{Z}$, denote by $\varrho_{B,e}:\overline{J}_{B}^{e}\to B$ the relative compactified Jacobian of degree $e$ associated with $\pi_{B}$. The fiber $\varrho_{B,e}^{-1}(b)$ over a point $b\in B$ is the moduli space of rank $1$, torsion-free sheaves of degree $e$ (or equivalently, with Euler characteristic $e+1-p_{a}({\mathcal{C}}_{b})$) on ${\mathcal{C}}_{b}$. Since $\pi_{B}$ admits, étale locally on $B$, a section which avoids the singularities of fibers of $\pi_{B}$, the smoothness of $\overline{J}_{B}^{e}$ is equivalent to that of $\overline{J}_{B}^{0}$. Hence, the following result follows directly from [10, Corollary B.3].
Proposition 3.5 ([10, Corollary B.3]).
If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a locally versal family of integral, locally planar curves, then $\overline{J}_{B}^{e}$ is smooth for any $e\in\mathbb{Z}$.